# Selective Population Protocols

### v2

@@ 4 Suitable Programming Environment
1def first_wins(X,Y,Z):
2  while X$>0$ and Y$>0$ in parallel:
3    Dec(X); Dec(Y); Inc(Inc(Z));
4  if X$>0$: return true
5  else: return false
6
7def rewrite(X,Y):
8  while X$>0$ in parallel:
9    Dec(X); Inc(Y);
10
11def M-protocol():
12  if first_wins($R$,$G$,$N$): // $R>0$
13    rewrite($R\vee N$,$R^{*}$);
14  else: // $G>0$
15    rewrite($G\vee N$,$G^{*}$);
@@ A.1 Fast multiplication
[…]
Stage 1: $|Y|\leftarrow|Y|/2$, remainder $b$ $(1a)\;$ $L_{in}+\mathcal{G}_{Y}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ L_{out}$ $(1b)\;$ $L_{in}+\mathcal{G}_{Y}|Y$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ L+A$ $(2a)\;$ $A+\mathcal{G}_{Y}|Y$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ A+A$ $(2b)\;$ $A+\mathcal{G}_{Y}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ A^{\prime}$ $(3a)\;$ $A^{\prime}+\mathcal{G}_{A}|A/A^{\prime}$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ B+free$ $(3b)\;$ $A^{\prime}+\mathcal{G}_{A}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ A^{\prime\prime}$ $(4a)\;$ $L+\mathcal{G}_{A}|A^{\prime\prime}$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ L^{1}+free$ $(4b)\;$ $L+\mathcal{G}_{A}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ L^{0}$ $(5)\;$ $B+\mathcal{G}_{A}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ Y$ $(6a)\;$ $L^{0}+\mathcal{G}_{B}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ L_{in}^{0}$ $(6b)\;$ $L^{1}+\mathcal{G}_{B}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ L_{in}^{1}$
     Stage 2: $|Z|\leftarrow|Z|+b|X|;|X|\leftarrow 2|X|$ $(1a)\;$ $L_{in}^{0}+\mathcal{G}_{X}|X$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ L^{\prime}+C^{0}$ $(1b)\;$ $L_{in}^{1}+\mathcal{G}_{X}|X$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ L^{\prime}+C^{1}$ $(2a)\;$ $C^{0}+\mathcal{G}_{X}|X$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ C^{0}+C^{0}$ $(2b)\;$ $C^{1}+\mathcal{G}_{X}|X$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ C^{1}+C^{1}$ $(2c)\;$ $C^{0}+\mathcal{G}_{X}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ C_{\star}^{0}$ $(2d)\;$ $C^{1}+\mathcal{G}_{X}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ C_{\star}^{1}$ $(3)\;$ $C_{\star}^{1}+\mathcal{G}_{f}|free$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ C_{\star}^{0}+Z$ $(4)\;$ $C_{\star}^{0}+\mathcal{G}_{f}|free$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ D+D$ $(5a)\;$ $D+\mathcal{G}_{C}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ X$ $(5b)\;$ $L^{\prime}+\mathcal{G}_{C}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ L^{\prime\prime}$ $(6)\;$ $L^{\prime\prime}+\mathcal{G}_{D}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ L_{in}$
@@ A.2 Quick-sort like selective ranking protocol
[…]
Phase 2: Establishing the rank of the pivot $P^{\prime}$ within the group $\mathcal{G}^{\prime}_{r}$, which is equal to $|B^{\prime}|+1$. The state value $v$ is used as a storage for partial results for the rank computation. $P^{\prime},A^{\prime},B^{\prime}\in\mathcal{G}^{\prime}_{r}$, $A^{\prime\prime},B^{\prime\prime},P^{\prime\prime}\in\mathcal{G}^{\prime\prime}_{r}$
($2a$, $A,A^{\prime}$ is rewritten for next phase as $A^{\prime\prime}$) $(A^{\prime}/B^{\prime}/P^{\prime},*,*)+\mathcal{G}^{\prime}_{r}|(A/A^{\prime},*,*)$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(A^{\prime}/B^{\prime}/P^{\prime},*,*)+(A^{\prime\prime},*,\_)$ ($2b$, counting $B^{\prime}$, agents already counted and rewritten to next phase) $(P^{\prime}/B^{\prime},*,x)+\mathcal{G}^{\prime}_{r}|(B^{\prime},*,y)$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(P^{\prime}/B^{\prime},*,x+y)+(B^{\prime\prime},*,\_)\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad$ ($2c$, pivot knows the group size, rewritten for the next phase) $(P^{\prime},*,x)+\mathcal{G}^{\prime}_{r}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(P^{\prime\prime},*,x)$
[…]
Phase 4: Randomizing new pivots for groups $\mathcal{G}^{\prime}_{r}$ (containing below agents) and $\mathcal{G}^{\prime}_{r+x+1}$ (containing above agents). Pivot’s values are set to their keys. Old pivot becomes a sole new agent in group $\mathcal{G}_{r+x}$ with type $R$. Computations proceed to Phase 1. $P^{\prime\prime\prime},P^{\prime\prime\prime\prime},P^{\prime\prime\prime\prime\prime},R\in\mathcal{G}_{r+x}$ $(4a,\text{all $N^{\prime}$ rewritten to $N$})\;$ $(P^{\prime\prime\prime},*,r)+\mathcal{G}^{\prime}_{r+x+1}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(P^{\prime\prime\prime\prime},*,r)$ $(4b,\text{indication of the new pivot in $\mathcal{G}_{r+x+1}$})\;$ $(P^{\prime\prime\prime\prime},*,r)+\mathcal{G}_{r+x+1}|(N,k,*)$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(P^{\prime\prime\prime\prime\prime},*,r)+(P[\mathcal{G}^{\prime}_{r+x+1}],k,k)$ $(4b,\text{in case no pivot candidates in $\mathcal{G}_{r+x+1}$})\;$ $(P^{\prime\prime\prime\prime},*,r)+\mathcal{G}_{r+x+1}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(P^{\prime\prime\prime\prime\prime},*,r)$ $(4c,\text{indication of the new pivot in $\mathcal{G}_{r}$})\;$ $(P^{\prime\prime\prime\prime\prime},*,r)+\mathcal{G}_{r}|(N,k,*)$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(R,*,*)+(P[\mathcal{G}^{\prime}_{r}],k,k)$ $(4c,\text{in case no pivot candidates in $\mathcal{G}_{r}$})\;$ $(P^{\prime\prime\prime\prime\prime},*,r)+\mathcal{G}_{r}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(R,*,*)$

### v3

@@ 4 Suitable Programming Environment
1def first_wins(X,Y,Z): 2 while X>0>0 and Y>0>0 in parallel: 3 Dec(X); Dec(Y); Inc(Inc(Z)); 4 if X>0>0: return true 5 else: return false 6 7def rewrite(X,Y): 8 while X>0>0 in parallel: 9 Dec(X); Inc(Y); 10 11def M-protocol(): 12 if first_wins($R$,$G$,NN): // $R>0$ 13 rewrite($\vee N,RR^{*}$); 14 else: // $G>0$ 15 rewrite($\vee N,GG^{*}$);
@@ 5 Final Comment
6 Acknowledgements
This research was supported in part by the National Science Centre, Poland, under project number 2020/39/B/ST6/03288 (Adam Gańczorz, Tomasz Jurdziński, Grzegorz Stachowiak).
This research was supported in part by the National Science Centre, Poland, under project number 2021/41/B/ST6/03691 (Jakub Kowalski).
@@ A.1 Fast multiplication
The multiplication protocol operates in $r=\lceil\log|Y|\rceil$ rounds which correspond to $r$ bits $b_{0},b_{1},\dots,b_{r-1}$ of binary expansion of the cardinality $|Y|.$ Each round is split into two stages. During Stage 1 the next bit $b_{i}$ of this expansion is computed starting from the least significant bit $b_{0}.$ This is done by testing parity of the current size of set $Y,$ which is reduced by half to the size |Y|div 2|Y|\ div\ 2 during each round. In Stage 2 set $Z$ is increased by $b_{i}\cdot 2^{i}|X|$ agents. Thus, after $j$ rounds the size of $Z$ is $\sum_{i=0}^{j}b_{i}\cdot 2^{i}|X|$, and after $r=\lceil\log|Y|\rceil$ rounds $\sum_{i=0}^{r}b_{i}\cdot 2^{i}|X|=|X|\cdot|Y|$.
Stage 1: $|Y|\leftarrow|Y|/2$, remainder $b$ $(1a)\;$ $L_{in}+\mathcal{G}_{Y}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ L_{out}\$$(1b)\;$ $L_{in}+\mathcal{G}_{Y}|Y$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ L+A$ $(2a)\;$ $A+\mathcal{G}_{Y}|Y$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ A+A$ $(2b)\;$ $A+\mathcal{G}_{Y}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ A^{\prime}\$$(3a)\;$ $A^{\prime}+\mathcal{G}_{A}|A/A^{\prime}$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ B+free$ $(3b)\;$ $A^{\prime}+\mathcal{G}_{A}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ A^{\prime\prime}\$$(4a)\;$ $L+\mathcal{G}_{A}|A^{\prime\prime}$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ L^{1}+free$ $(4b)\;$ $L+\mathcal{G}_{A}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ L^{0}\$$(5)\;$ $B+\mathcal{G}_{A}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ Y\$$(6a)\;$ $L^{0}+\mathcal{G}_{B}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ L_{in}^{0}\$$(6b)\;$ $L^{1}+\mathcal{G}_{B}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ L_{in}^{1}\$   Stage 2: $|Z|\leftarrow|Z|+b|X|;|X|\leftarrow 2|X|$ $(1a)\;$ $L_{in}^{0}+\mathcal{G}_{X}|X$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ L^{\prime}+C^{0}$ $(1b)\;$ $L_{in}^{1}+\mathcal{G}_{X}|X$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ L^{\prime}+C^{1}$ $(2a)\;$ $C^{0}+\mathcal{G}_{X}|X$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ C^{0}+C^{0}$ $(2b)\;$ $C^{1}+\mathcal{G}_{X}|X$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ C^{1}+C^{1}$ $(2c)\;$ $C^{0}+\mathcal{G}_{X}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ C_{\star}^{0}\$$(2d)\;$ $C^{1}+\mathcal{G}_{X}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ C_{\star}^{1}\$$(3)\;$ $C_{\star}^{1}+\mathcal{G}_{f}|free$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ C_{\star}^{0}+Z$ $(4)\;$ $C_{\star}^{0}+\mathcal{G}_{f}|free$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ D+D$ $(5a)\;$ $D+\mathcal{G}_{C}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ X\$$(5b)\;$ $L^{\prime}+\mathcal{G}_{C}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ L^{\prime\prime}\$$(6)\;$ $L^{\prime\prime}+\mathcal{G}_{D}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $\ L_{in}\$
@@ A.2 Quick-sort like selective ranking protocol
In groups $\mathcal{G}_{r}$, $\mathcal{G}^{\prime}_{r}$, $\mathcal{G}^{\prime\prime}_{r}$, the index $r$ is the minimal rank of an element within a group. Agent states are encoded as triplets $(t,k,v)$, where $t$ is the type of the agent, used to differentiate its role in the computation process, $k$ is its key, constant for each agent, and $v$ encodes additional value required for computations, its meaning depends on the agent type and phase of the protocol. If, inside a transition rule, an element from state’s tuple remains unchanged, we mark such situation using *. When the value does not matter for further computations _\_ is used instead. If it is not clear from the context of the rule to which group an agent of type $T$ belongs, we write it explicitly as $T[\mathcal{G}]$. Phase 0: Choosing the initial pivot for the entire population of agents. All agents start as $(L,k,\_)\in\mathcal{G}^{\prime}_{1}$, where $k$ is each agent’s key, know to them and used for comparion with other keys. $N\in\mathcal{G}_{1}$, $L,P\in\mathcal{G}^{\prime}_{1}$ $(0a,\text{ leader election})\;$ $(L,*,*)+\mathcal{G}^{\prime}_{1}|(L,*,*)$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(L,*,*)+(N,*,\_)\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt$ $(0b,\text{leader becomes pivot for the next phase})\;$ $(L,k,*)+\mathcal{G}^{\prime}_{1}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(P,k,k)\$Phase 1: For every group $\mathcal{G}_{r}$ initially containing only neutral agents NN, and $\mathcal{G}^{\prime}_{r}$ with a single pivot $P$, the $\mathcal{G}_{r}$ members are transferred to $\mathcal{G}^{\prime}_{r}$, divided into being above (AA) or below ($B$) the pivot, and learn the key of the pivot (stored in the state’s third value $v$). $N\in\mathcal{G}_{r}$, $P,P^{\prime},A,A^{\prime},B,B^{\prime}\in\mathcal{G}^{\prime}_{r}$ $(1a,\text{if }k<l\text{: pivot sets neutral as above})\;$ $(P,k,k)+\mathcal{G}_{r}|(N,l,\_)$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(P,k,k)+(A,l,k)\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt$ $(1a,\text{if }k>l\text{: pivot sets neutral as below})\;$ $(P,k,k)+\mathcal{G}_{r}|(N,l,\_)$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(P,k,k)+(B,l,k)$ $(1b,\text{if }k<l\text{: above/below sets neutral as above})\;$ $(A/B,*,k)+\mathcal{G}_{r}|(N,l,\_)$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(A/B,*,k)+(A,l,k)$ $(1b,\text{if }k>l\text{: above/below sets neutral as below})\;$ $(A/B,*,k)+\mathcal{G}_{r}|(N,l,\_)$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(A/B,*,k)+(B,l,k)$ $(1c,\text{pivot/above rewritten for the next phase})\;$ $(P/A,*,*)+\mathcal{G}_{r}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(P^{\prime}/A^{\prime},*,0)\$$(1d,\text{below rewritten for the next phase})\;$ $(B,*,*)+\mathcal{G}_{r}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(B^{\prime},*,1)\$Phase 2: Establishing the rank of the pivot $P^{\prime}$ within the group $\mathcal{G}^{\prime}_{r}$, which is equal to $|B^{\prime}|+1$. The state value $v$ is used as a storage for partial results for the rank computation. $P^{\prime},A^{\prime},B^{\prime}\in\mathcal{G}^{\prime}_{r}$, $A^{\prime\prime},B^{\prime\prime},P^{\prime\prime}\in\mathcal{G}^{\prime\prime}_{r}$ ($2a$, $A,A^{\prime}$ is rewritten for next phase as $A^{\prime\prime}$) $(A^{\prime}/B^{\prime}/P^{\prime},*,*)+\mathcal{G}^{\prime}_{r}|(A/A^{\prime},*,*)$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(A^{\prime}/B^{\prime}/P^{\prime},*,*)+(A^{\prime\prime},*,\_)$ ($2b$, counting $B^{\prime}$, agents already counted and rewritten to next phase) $(P^{\prime}/B^{\prime},*,x)+\mathcal{G}^{\prime}_{r}|(B^{\prime},*,y)$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(P^{\prime}/B^{\prime},*,x+y)+(B^{\prime\prime},*,\_)\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt$ ($2c$, pivot knows the group size, rewritten for the next phase) $(P^{\prime},*,x)+\mathcal{G}^{\prime}_{r}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(P^{\prime\prime},*,x)\$Phase 3: Resetting agent, above and below lose their type and become neutrals in proper groups, either $\mathcal{G}_{r}$ or $\mathcal{G}_{r+x+1}$. Additionally, above agents learn the pivot’s rank $r+x$. State value is used for storing index of the agent’s old group or its rank within it, depending on particular need. $B^{\prime\prime},P^{\prime\prime}\in\mathcal{G}^{\prime\prime}_{r}$, $N^{\prime}\in\mathcal{G}^{\prime}_{r+x+1}$ $(3a,\text{below become neutral in group $\mathcal{G}_{r}$ })\;$ $(B^{\prime\prime},*,*)+\mathcal{G}^{\prime}_{r}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(N[\mathcal{G}_{r}],*,\_)\$$(3b,\text{above become neutral $N^{\prime}\in\mathcal{G}^{\prime}_{r+x+1}$})\;$ $(P^{\prime\prime},*,x)+\mathcal{G}^{\prime\prime}_{r}|(A^{\prime\prime},*,*)$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(P^{\prime\prime},*,x)+(N^{\prime},*,r)$ $(3b,\text{above become neutral $N^{\prime}\in\mathcal{G}^{\prime}_{r+x+1}$})\;$ $(N^{\prime},*,r)+\mathcal{G}^{\prime\prime}_{r}|(A^{\prime\prime},*,*)$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(N^{\prime},*,r)+(N^{\prime},*,r)$ $(3c,\text{$N^{\prime}$ become standard neutral in $\mathcal{G}_{r+x+1}$ })\;$ $(N^{\prime},*,r)+\mathcal{G}^{\prime\prime}_{r}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(N[\mathcal{G}_{r+x+1}],*,\_)\$$(3d,\text{pivot moves to group corresponding to its rank})\;$ $(P^{\prime\prime},*,x)+\mathcal{G}^{\prime\prime}_{r}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(P^{\prime\prime\prime}[\mathcal{G}_{r+x}],*,r)\$Phase 4: Randomizing new pivots for groups $\mathcal{G}^{\prime}_{r}$ (containing below agents) and $\mathcal{G}^{\prime}_{r+x+1}$ (containing above agents). Pivot’s values are set to their keys. Old pivot becomes a sole new agent in group $\mathcal{G}_{r+x}$ with type $R$. Computations proceed to Phase 1. $P^{\prime\prime\prime},P^{\prime\prime\prime\prime},P^{\prime\prime\prime\prime\prime},R\in\mathcal{G}_{r+x}$ $(4a,\text{all $N^{\prime}$ rewritten to $N$})\;$ $(P^{\prime\prime\prime},*,r)+\mathcal{G}^{\prime}_{r+x+1}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(P^{\prime\prime\prime\prime},*,r)\$$(4b,\text{indication of the new pivot in $\mathcal{G}_{r+x+1}$})\;$ $(P^{\prime\prime\prime\prime},*,r)+\mathcal{G}_{r+x+1}|(N,k,*)$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(P^{\prime\prime\prime\prime\prime},*,r)+(P[\mathcal{G}^{\prime}_{r+x+1}],k,k)$ $(4b,\text{in case no pivot candidates in $\mathcal{G}_{r+x+1}$})\;$ $(P^{\prime\prime\prime\prime},*,r)+\mathcal{G}_{r+x+1}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(P^{\prime\prime\prime\prime\prime},*,r)\$$(4c,\text{indication of the new pivot in $\mathcal{G}_{r}$})\;$ $(P^{\prime\prime\prime\prime\prime},*,r)+\mathcal{G}_{r}|(N,k,*)$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(R,*,*)+(P[\mathcal{G}^{\prime}_{r}],k,k)$ $(4c,\text{in case no pivot candidates in $\mathcal{G}_{r}$})\;$ $(P^{\prime\prime\prime\prime\prime},*,r)+\mathcal{G}_{r}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $(R,*,*)\$